# ExploraCoder: Advancing code generation for multiple unseen APIs via planning and chained exploration

### v1

@@ A.1 Evaluating API knowledge retention in API-pretrained and API-untrained base models
ExploraCoder uses some widely-adopted design elements/paradigms (such as CoT, RAG, code execution), which may also be seen in some recently popular Swe-bench agents (Yang et al., 2024; Zhang et al., 2024; Ma et al., 2024a). However, overlapping in framework elements does not represent resemblance in ?. We highlight below the key distinctions in application focus and methodology that clearly differentiate our work. And we further discuss the novelty and improvements
A.3 Difference from Swe-Bench Code Agents
In this section, we address the differences between our framework, ExploraCoder, and the recently popular Swe-Bench code agents. While there may be some similarities in certain elements, we highlight key distinctions in application focus and methodology that clearly differentiate our work.
Shared Elements. ExploraCoder shares certain widely adopted elements/paradigms (such as planning, CoT reasoning, RAG, and execution) with Swe-Agent (Yang et al., 2024), AutoCodeRover (Zhang et al., 2024), and other advanced Swe-Bench agents (Liu et al., 2024b; Ma et al., 2024a). However, note that these paradigms are commonly used in recent research (Le et al., 2024; Ni et al., 2024), and overlapping adoption of elements does not suggest significant resemblance in overall framework design. The novelty and improvements in ExploraCoder’s framework designs are further discussed in Appendix A.4.
Scenario Differences. The application focus of ExploraCoder is fundamentally different from that of Swe-Bench agents. Swe-Bench agents primarily rely on a small, predefined set of repository tools to identify buggy lines and edit existing code. In contrast, ExploraCoder synthesizes complex, unseen API invocation code. Using the terminology of agentic systems, our framework addresses the challenge of selecting appropriate tools from a large, third-party toolset and ensuring their correct usage—a scenario that falls outside the scope of Swe-Bench agents, and many other agentic systems. Therefore, these agents are not directly applicable to our scenario of library-oriented code generation.
Methodological Differences. More importantly, ExploraCoder employs a distinctive methodology. Firstly, ExploraCoder adopts a non-agent approach. Xia et al. (2024) summarizes existing agent systems as adhering to a “reason-then-act” framework, iteratively planning future actions based on transparent action space. Such frameworks require additional heavily annotated tool use descriptions and often lead to uncontrollable decision and execution. ExploraCoder, however, utilizes a document-enhanced planner specifically designed to accommodate domain-specific planning logic. This planner generates an integrated set of subtasks in the absence of full API list, providing more controllable and interpretable iteration steps, as detailed in Section 3.2. Moreover, ExploraCoder differs from most code agents during code generation phase by focusing on partial code generation in each intermediate step. Whereas code agents typically generate complete code and refine it iteratively in their subsequent steps–a process similar to our ”CAPIR + Self-Repair” baseline. Experimental results shows the superiority of our approach. These distinct designs set ExploraCoder apart in terms of both framework pipeline and methodology.
A.4 Additional Discussion of Novelty in Methodology
(1) A novel iterative generation method: We designed a novel Chain-of-API-Exploration (CoAE) mechanism for step-wise code generation. In each intermediate step, CoAE generates partial API invocations based on limited API documentation and leverages executability signals to rectify coding decisions in real time.
The key difference between CoAE and other (CoT-inspired and execution-driven) iterative code generation methods, such as CodeChain (Le et al., 2024), Self-Debug (Chen et al., 2023), and Swe-Agent (Yang et al., 2024), lies in that it focuses on each intermediate subtask within a targeted integral code solution and uses execution as step-wise reward signal. Specifically, Self-Debug and Swe-Agent applies end-to-end code/patch generation and iteratively edit the code/patch based on new information from executing the code/reproduction script. CodeChain generates modularized code snippets but does not consider step-wise execution signals and fails to effectively utilize the logical relationships between modules. Additionaly, CodeChain is designed for self-contained code generation thus cannot be applied to coding task with external API invocations.
(2) An Exploratory-Programming-inspired Framework: We designed a unified framework, ExploraCoder, to integrate CoAE with other effective techniques. ExploraCoder is inspired by the exploratory programming paradigm observed in human developers (Beth Kery & Myers, 2017), who actively read API documents and gather more API usage experience from trial execution.
While ExploraCoder integrates existing techniques such as CoT reasoning, RAG, and execution-then-debug, its framework design introduces non-trivial innovations beyond merely applying these techniques to the domain of unseen API usage. Specifically:
  • Integration of CoT reasoning: Unlike the trivial decomposition of subtasks as CoT instructions, ExploraCoder enhances CoT reasoning by utilizing the exploration trace produced by CoAE as enriched CoT instructions.
  • Debugging mechanisms: ExploraCoder incorporates debugging at each CoAE intermediate step, rather than debugging the final solution as done in CAPIR+Self-Repair or many Swe-Bench agents.
  • API retrieval: Beyond retrieving API docs based solely on semantic similarity, we refine the recommended API set by dynamically parsing the CoAE steps to identify APIs that are more relevant and usable.
Experimental results also indicates the superiority of ExploraCoder’s design choice compared to the aforementioned trivial technique application.
A.5 Quantitative analysis analysis for CoAE
ExploraCoder leverages API invocation experience from CoAE to enhance the quality of final solution generation. Intuitively, the quality of exploration subtasks within CoAE is closely related to the quality of the final solutions. We first examine the overall success rate of CoAE subtasks and final solutions in ExploraCoder in Table 8. We observe that ExploraCoder’s performance declines on Torchdata-Manual that involves longer exploration chains, with a 17.48% lower CoAE subtask success rate and a 19.83% lower final solution success rate comparing to Torchdata-Github. This could be attributed to failures in the intermediate steps blocking the fully correctness of the final solutions.
Table 8: We compare the success rate of ExploraCoder on two datasets.
Base model Dataset length of CoAE CoAE Success Solution Success
GPT-3.5-turbo Torchdata-Github 4.06 63.00% 57.30%
Torchdata-Manual 8.22 42.19% 33.81%
GPT-4-0613 Torchdata-Github 5.28 71.74% 61.87%
Torchdata-Manual 8.80 57.60% 45.70%
To further explore this relationship, we conducted a quantitative analysis, examining how the number of CoAE subtasks and their success rates affect the pass rate and overall success rate of the final solutions. We illustrate the correlation in a scatter plot (Figure 2), using results from the best-performing base model, GPT-4-0613, on our Torchdata-Manual benchmark. This dataset was chosen due to its diverse range of API invocation complexities, which result in varied numbers of decomposed subtasks, providing better visualization of the relationship. From Figure 2(a) to 2(d), we observe that both the pass rate and success rate of the final solutions positively correlate with the CoAE subtask success rate. Subtasks with higher success rates, particularly those with a success rate of 1, are more likely to generate successful or passing final solutions. Interestingly, the number of subtasks doesn’t appear to have a significant direct impact. However, as shown in Figures 2(a) and 2(b), without self-debugging, problems with a higher subtask number (ranging from 10 to 13) tend to have lower subtask success rate as the subtask number increase. This may be due to the increased complexity of inter-task API interactions, which can overwhelm LLMs. When the self-debug mechanism is introduced in ExploraCoder*, we observe in Figures 2(c) and 2(d) a notable improvement in the overall subtask success rate, even for cases with higher subtask numbers. This leads to more successful and passing final solutions. The improvement can be attributed to ExploraCoder’s ability to correct typos and simple API interaction errors in each subtask, thereby gaining richer API usage experience and exploiting it to the final solution generation.
(a) success-fail division
for final solution in
ExploraCoder
(b) pass-fail division
for final solution in
ExploraCoder
(c) success-fail division
for final solution in
ExploraCoder*
(d) pass-fail division
for final solution in
ExploraCoder*
A.6 The effectiveness of task planning module in ExploraCoder
#API invocation #decomposed subtask API per subtask
Torchdata-Github 4.26 4.06 1.04
Torchdata-Manual 9.94 8.22 1.21
Table 9: Summary of decomposed subtask statistics.
A.7 The effectiveness of API recommendation module in ExploraCoder
Table 10: Choices of hyperparameters for ExploraCoder’s Retrival module: we tested retrieval modules effectiveness by 3 aspects: choice of embedding model, retrieval index(Desc: API description, Path: API import path, Desc*: truncated first sentense of API description), retrieval method(ST: single-task, MT: multi-task).
Retreival
Model
Retrieval
Index
Retrieval
Method
Racall@3 Recall@5 Recall@10 Revall@15
BM25 Desc ST 10.23 13.02 18.07 23.12
BGE Desc ST 12.57 16.50 25.49 32.81
ADA Desc ST 13.24 17.88 28.83 36.18
ADA Path ST 10.91 13.46 16.48 19.75
ADA Desc* ST 16.00 22.40 32.66 40.21
ADA Path+Desc* ST 19.64 24.66 34.28 40.91
ADA Path+Desc* MT 24.89 31.25 43.63 52.07
We evaluate the effectiveness of our API recommendation module using the Torchdata-AR benchmark (Ma et al., 2024b). Our experiments explored the performance difference in variant hyperparameters of the retrieval model, retrieval index, and retrieval methodology. Specifically, we assess the performance of a lexical method BM25 and two SOTA dense retrieval models, bge-large-en-v1.5, and text-embedding-ada-002. Our results show that dense retrieval models significantly outperform BM25, and we choose the best-performing text-embedding-ada-002 as our retreival model in our experiments.
For retrieval index construction, we observe that leveraging semantic information from both the API import paths and the first sentence of API descriptions yields the best performance. Notably, using only the first sentence of the API description outperforms using the entire description. A possible explanation is that the first sentence typically provides a concise summary of the API, which is sufficient for retrieval purposes. In contrast, the remaining content often introduces more detailed but potentially distracting information, such as parameter details and API behavior. In our experiment, we construct retrieval index by concating the API import path and the first line of API descriptions.
We also evaluate the impact of multi-task retrieval (MT), where complex problems are decomposed into multiple subtasks, each retrieving its own relevant APIs. The retrieval results are then reranked across subtasks, and the top-k APIs are selected. Our findings indicate that MT retrieval significantly improves recall compared to single-task retrieval (ST), where the complex problem is treated as a single task to retrieve APIs.
A.8 Construction of Torchdata-Manual
The Torchdata-Manual benchmark is designed to provide complex programming problems that require the use of multiple Torchdata APIs. It follows the style of prior unseen library benchamarks (Zan et al., 2023; Ma et al., 2024b) , consisting of a natural language task description, code context, canonical solutions, and test cases. The construction process is outlined as follows:
Expert Review. Finally, we invited two Python programmers, each with four years of experience, to review the dataset and suggest adjustments. Specifically, we ask the experts to examine on 4 aspect of the crafted programming problems (1) The executability of the canonical solution, (2) The intuitiveness of the API usage, (3) The rigor of the test cases, (3) The meaningfulness of the task requirements. If any issues were identified in these aspects, the experts discussed them with the task creators and revised the tasks accordingly. This step ensured the overall quality and correctness of the benchmark.
A.9 Additional implementation details
Torchdata is a library that facilitate multiple data processing operations. For task planning module, we ask GPT-3.5-turbo-0125 (API-untrained model) to summarize Torchdata’s purpose, key concepts, and API division logic based on Torchdata’s README page [^8] https://github.com/pytorch/data/blob/v0.7.1/README.md. The summarized results are presented in Listing 2. We also extracted few-shot API invocation planners demonstrated in Listing 3 following Ma et al. (2024b)’s approach. And both information are used for invocation task planning. Unlike the detailed functionalities for each APIs, the summarization and planners demonstrations give high-level insights into the library, facilitating better planning and reasoning for LLMs (Zheng et al., 2024). We use such summarization to represent limited domain knowledge for task planning, and no further detailed API usage information is leaked for problem solving. We also demonstrate ExploraCoder’s prompts in Listing 4 - 7.
@@ 3. Iterate over the datapipe and output both the URL and the first line of the text file
A.10 Case study
@@ 6. Cache and hold the processed data in memory
Listing 10: A failed example for Self-Repair.
[requirement]
--------- initial candidate ---------
[code]
from torchdata.datapipes.iter import *
from typing import List, Tuple
import torch
file_name = paragraph[0]
paragraph_content = paragraph[1]
return paragraph_content
def build_text_data_pipe(
batch_size: int=16, # Batch size
data_pipe = data_pipe >> Shuffler()
return data_pipe
list_dp = list(build_text_data_pipe(root="./torchdata-programming-tasks/task_82", batch_size=4))
[execution result]
Status: failed
Error message:
TypeError: __init__() missing 1 required positional argument: root
Error traceback at:
- line 23 in <module>: list_dp = list(build_text_data_pipe(root="./torchdata-programming-tasks/task_82", batch_size=4))
--------- round 1 self-repair ---------
[feedback]
[code]
from torchdata.datapipes.iter import *
from typing import List, Tuple
import torch
file_name = paragraph[0]
paragraph_content = paragraph[1]
return paragraph_content
def build_text_data_pipe(
batch_size: int=16, # Batch size
data_pipe = data_pipe >> Shuffler()
return data_pipe
list_dp = list(build_text_data_pipe(root="./torchdata-programming-tasks/task_82", batch_size=4))
Status: failed
Error message:
TypeError: __init__() missing 1 required positional argument: root
Error traceback at:
- line 23 in <module>: list_dp = list(build_text_data_pipe(root="./torchdata-programming-tasks/task_82", batch_size=4))
--------- round 2 self-repair ---------
[feedback]
[code]
from torchdata.datapipes.iter import *
from typing import List, Tuple
import torch
file_name = paragraph[0]
paragraph_content = paragraph[1]
return paragraph_content
def build_text_data_pipe(
batch_size: int=16, # Batch size
data_pipe = data_pipe >> Shuffler()
return data_pipe
list_dp = list(build_text_data_pipe(root="./torchdata-programming-tasks/task_82", batch_size=4))
[execution result]
Status: failed
Error message:
TypeError: __init__() missing 1 required positional argument: root
Error traceback at:
- line 23 in <module>: list_dp = list(build_text_data_pipe(root="./torchdata-programming-tasks/task_82", batch_size=4))
--------- round 3 self-repair ---------
[feedback]
[code]
from torchdata.datapipes.iter import *
from typing import List, Tuple
import torch
file_name = paragraph[0]
paragraph_content = paragraph[1]
return paragraph_content
def build_text_data_pipe(
batch_size: int = 16, # Batch size
data_pipe = data_pipe >> Shuffler()
return data_pipe
list_dp = list(build_text_data_pipe(root="./torchdata-programming-tasks/task_82", batch_size=4))
--------- round 4-5 self-repair ---------
...omit for simplicity...
--------- round 6 self-repair ---------
[feedback]
[code]
from torchdata.datapipes.iter import *
from typing import List, Tuple
import torch
file_name = paragraph[0]
paragraph_content = paragraph[1]
return paragraph_content
def build_text_data_pipe(
batch_size: int = 16, # Batch size
data_pipe = data_pipe >> Shuffler()
return data_pipe
list_dp = list(build_text_data_pipe(root="./torchdata-programming-tasks/task_82", batch_size=4))
[execution result]
Status: failed
Error message:
TypeError: __init__() missing 1 required positional argument: root
Error traceback at:
- line 23 in <module>: list_dp = list(build_text_data_pipe(root="./torchdata-programming-tasks/task_82", batch_size=4))
--------- Analysis ---------
Listing 11: A successful example for self-debug on ExploraCoder.
@@
--------- failed experimental code ---------
@@
This exception is thrown by __iter__ of FileOpenerIterDataPipe(datapipe=ZipArchiveLoaderIterDataPipe, encoding=None, length=-1, mode=’r’)
@@
--------- fixed by self-debug ---------
@@
Text Data from torchdata-programming-tasks/task_89/file_0.zip/file_0_0.txt: b’...’

### v2

@@ A.6 Evaluating prior API knowledge retention in API-pretrained and API-untrained base models
A.8 Additional implementation details
Torchdata is a library that facilitate multiple data processing operations. For task planning module, we ask GPT-3.5-turbo-0125 (API-untrained model) to summarize Torchdata’s purpose, key concepts, and API division logic based on Torchdata’s README page [^5] https://github.com/pytorch/data/blob/v0.7.1/README.md. The summarized results are presented in Listing 2. We also extracted few-shot API invocation planners demonstrated in Listing 3 following Ma et al. (2024)’s approach. And both information are used for invocation task planning. Unlike the detailed functionalities for each APIs, the summarization and planners demonstrations give high-level insights into the library, facilitating better planning and reasoning for LLMs (Zheng et al., 2024). We use such summarization to represent limited domain knowledge for task planning, and no further detailed API usage information is leaked for problem solving. We also demonstrate ExploraCoder’s prompts in Listing 4 - 7.
@@ 3. Iterate over the datapipe and output both the URL and the first line of the text file
A.9 Case study
We have conducted a series of case studies, here we provide examples of different methods (naive RAG, ExploraCoder, Self-Repair) solving the same example problem from our benchmark. We also provided a case study of ExploraCoder*, where we demonstrate the self-debug trace at an failed intermediate subtask in CoAE. For each example, we provide discussion and analysis in the end of the listing.
@@ 6. Cache and hold the processed data in memory
Listing 10: A successful example for intermediate self-debug on ExploraCoder.
@@
----------------- failed experimental code -----------------
@@
This exception is thrown by __iter__ of FileOpenerIterDataPipe(datapipe=
ZipArchiveLoaderIterDataPipe, encoding=None, length=-1, mode=’r’)
@@
------------------- fixed by self-debug -------------------
@@
Text Data from torchdata-programming-tasks/task_89/
file_0.zip/file_0_0.txt: b’...’
@@
Listing 11: A failed example for Self-Repair.
[requirement]
-------------------- initial candidate --------------------
[code]
from torchdata.datapipes.iter import *
from typing import List, Tuple
import torch
file_name = paragraph[0]
paragraph_content = paragraph[1]
return paragraph_content
def build_text_data_pipe(
batch_size: int=16, # Batch size
data_pipe = data_pipe >> Shuffler()
return data_pipe
list_dp = list(build_text_data_pipe(root="./torchdata-programming
-tasks/task_82", batch_size=4))
[execution result]
Status: failed
Error message:
TypeError: __init__() missing 1 required positional argument: root
Error traceback at:
- line 23 in <module>: list_dp = list(build_text_data_pipe(root="./torchdata-programming
-tasks/task_82", batch_size=4))
------------------- round 1 self-repair -------------------
[feedback]
[code]
from torchdata.datapipes.iter import *
from typing import List, Tuple
import torch
file_name = paragraph[0]
paragraph_content = paragraph[1]
return paragraph_content
def build_text_data_pipe(
batch_size: int=16, # Batch size
data_pipe = data_pipe >> Shuffler()
return data_pipe
list_dp = list(build_text_data_pipe
(root="./torchdata-programming-tasks/task_82", batch_size=4))
Status: failed
Error message:
TypeError: __init__() missing 1 required positional argument: root
Error traceback at:
- line 23 in <module>: list_dp = list
(build_text_data_pipe(root="./torchdata-programming-
tasks/task_82", batch_size=4))
------------------- round 2 self-repair -------------------
[feedback]
[code]
from torchdata.datapipes.iter import *
from typing import List, Tuple
import torch
file_name = paragraph[0]
paragraph_content = paragraph[1]
return paragraph_content
def build_text_data_pipe(
batch_size: int=16, # Batch size
data_pipe = data_pipe >> Shuffler()
return data_pipe
list_dp = list(build_text_data_pipe(root="./torchdata
-programming-tasks/task_82", batch_size=4))
[execution result]
Status: failed
Error message:
TypeError: __init__() missing 1 required positional argument: root
Error traceback at:
- line 23 in <module>: list_dp = list(build_text_data_pipe(root="./torchdata-programming
-tasks/task_82", batch_size=4))
------------------- round 3 self-repair -------------------
[feedback]
Listing 12: A failed example for Self-Repair (Cont.).
[code]
from torchdata.datapipes.iter import *
from typing import List, Tuple
import torch
file_name = paragraph[0]
paragraph_content = paragraph[1]
return paragraph_content
def build_text_data_pipe(
batch_size: int = 16, # Batch size
data_pipe = data_pipe >> Shuffler()
return data_pipe
list_dp = list(build_text_data_pipe(
root="./torchdata-programming-tasks/task_82", batch_size=4))
------------------ round 4-5 self-repair ------------------
...omit for simplicity...
------------------- round 6 self-repair -------------------
[feedback]
[code]
from torchdata.datapipes.iter import *
from typing import List, Tuple
import torch
file_name = paragraph[0]
paragraph_content = paragraph[1]
return paragraph_content
def build_text_data_pipe(
batch_size: int = 16, # Batch size
data_pipe = data_pipe >> Shuffler()
return data_pipe
list_dp = list(build_text_data_pipe(
root="./torchdata-programming-tasks/task_82", batch_size=4))
[execution result]
Status: failed
Error message:
TypeError: __init__() missing 1 required positional argument: root
Error traceback at:
- line 23 in <module>: list_dp = list(build_text_data_pipe(
root="./torchdata-programming-tasks/task_82", batch_size=4))
------------------------- Analysis -------------------------